OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

1. Find the slope. 2. Use the point-slope formula to find the equation of the line. 3. Set y equal to 0 to find the x intercept.

OpenStudy (anonymous):

Can u show me how to do it plz

OpenStudy (anonymous):

I can give you the formulas... http://www.purplemath.com/modules/slope.htm slope = \(\dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

Plug in your x and y values.

OpenStudy (anonymous):

3-(-3)/-5-1?

OpenStudy (anonymous):

Yes, now do the arithmetic to get the slope.

OpenStudy (anonymous):

6/-6

OpenStudy (anonymous):

I'm not sure what I'm doing

OpenStudy (anonymous):

Simplify your fraction.

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

6 divided by -6 = what? Point-Slope Formula: \(y – y1 = m(x – x1)\)

OpenStudy (anonymous):

When multiplying or dividing an odd number of negatives, the answer is negative.

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Now what shall I do

OpenStudy (anonymous):

Purple Math, the best resource for high school math questions.... http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (anonymous):

Now, plug in the values for \(m,y_1,x_1\)

OpenStudy (anonymous):

m = slope

OpenStudy (anonymous):

I get it thanks

OpenStudy (anonymous):

Your Welcome.